Hello. Can you help me solve this word problem please? Thank you:
Half the books you own are mysteries. One third are nonfiction. The other four books are science fiction. How many oy your books are mysteries?

Answer by rfer(16322)

1/2+1/3=
3/6+2/6=5/6
------------
4/(1/6)=
4*6/1=24 total
--------------
24*1/2=24/2=12 mystery